• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 01:34:05 on April 07, 2008

    Entry number 230474

    keyword=Rear chamber problem summary

    If you've been following the Halog you've seen that we had a problem with the rear wirechamber in the Hadron stack tonight. (Persistent overcurrent trips on ramping the rear-most field planes.)

    Xin and I did some troubleshooting and it looks like one of the field wires is coupled to a signal wire on the U13 connector. Pulling that card allowed us to bring the chamber back up. Fortunately the wires on U13 run diagonally across the lower-left corner of the chamber from the target perspective (lower-right if you're looking at the stack from behind). I can't imagine the loss of those wires will significantly impact the solid angle coverage -- I think we got lucky. If nothing else goes wrong I expect we can run out the experiment without swapping chambers.

    When we came in this evening it was noted that the ramp rate for the chamber HV had been changed from 10 V/s to 500 V/s. This change occured sometime between 8pm Saturday night and 10pm Sunday. It is not clear if a shift worker made the change, or if the HV control software had an initialization problem at some point and loaded a high default(?) ramp rate. We put a sign up on the computer but a software-based monitor would be a better solution.